# Bilinear MLPs enable weight-based mechanistic interpretability

### v1

@@ 1 Introduction
One path to better circuit discovery is to use more inherently interpretable architectures. Previous work constrains the model into using human-understandable components or concepts (Koh et al., 2020; Chen et al., 2019), but this typically requires labeled training data for the predefined concepts and involves a trade-off in accuracy compared to learning the best concepts for performance (Henighan et al., 2023). Ideally, we could more easily extract and interpret the concepts that models naturally learn rather than force the model to use particular concepts. To this end, Sharkey (2023) suggested that bilinear layers of the form $g({\bm{x}})=({\bm{W}}{\bm{x}})\odot({\bm{V}}{\bm{x}})$ are intrinsically interpretable because their computations can be expressed in terms of linear operations with a third order tensor. This enables the use of tensor or matrix decompositions to understand the weights. Moreover, bilinear layers outperform ReLU-based transformers in language modeling (Shazeer, 2020) and have performance only slightly below SwiGLU, which is prevalent in competitive models today (Touvron et al., 2023).
We show that bilinear MLPs can be decomposed into functionally relevant, interpretable components by directly decomposing the weights, without using inputs. These decompositions reveal a low-rank structure in bilinear MLPs trained across various tasks. In summary, this paper demonstrates that bilinear MLPs are an interpretable drop-in replacement for ordinary MLPs in a wide range of settings. Our contributions are as follows:
@@ 2 Background
Throughout, we use conventional notation as in Goodfellow et al. (2016). Scalars are denoted by $s$, vectors by $\bm{v}$}, matrices by $\bm{M}$}, and third-order tensors by $\bm{\mathsfit{T}}$}. The entry in row $i$ and column $j$ of a matrix $\bm{M}$} is a scalar and therefore denoted as $m_{ij}$. We denote taking row $i$ or column $j$ of a matrix by ${\bm{m}}_{i:}$ and ${\bm{m}}_{:j}$ respectively. We use $\odot$ to denote an element-wise product and $\cdot_{\text{axis}}$ to denote a product of tensors along the specified axis.
Defining bilinear MLPs. Modern Transformers (Touvron et al., 2023) feature Gated Linear Units (GLUs), which offer a performance gain over standard MLPs for the same number of parameters (Shazeer, 2020). GLU activations consist of the component-wise product of two linear up-projections, $\bm{W}$} and $\bm{V}$}, one of which is passed through a nonlinear activation function $\sigma$(Equation 1). We omit biases for brevity.
@@
A bilinear layer is a GLU variant that omits the nonlinear activation function $\sigma$. Bilinear layers beat ordinary ReLU MLPs, and perform almost as well as SwiGLU on language modeling tasks (Shazeer, 2020). We corroborate these findings in Appendix G, and show that bilinear layers achieve equal loss when keeping training time constant and marginally worse loss when keeping data constant.
@@
We call the matrix ${\bm{w}}_{a:}{\bm{v}}_{a:}^{T}={\bm{B}}_{a::}$ an interaction matrix since it defines how each pair of inputs interact for a given output dimension $a$.
The collection of interaction matrices across the output axis can be organized into the third-order bilinear tensor, $\bm{\mathsfit{B}}$}, with elements $b_{aij}=w_{ai}v_{aj}$, illustrated in Figure 1A. The bilinear tensor allows us to easily find the interaction matrix for a specific output direction $\bm{u}$} of interest by taking a product along the output axis, ${\bm{u}}\ \cdot_{\text{out}}{\bm{\mathsfit{B}}}$, equal to a weighted sum over the neuron-basis interaction matrices, $\sum_{a}u_{a}{\bm{w}}_{a:}{\bm{v}}_{a:}^{T}$.
Simplifications due to symmetry. Because an interaction matrix is always evaluated with two copies of the input $\bm{x}$}, it contains redundant information that does not contribute to the activation of a layer. Specifically, any square matrix can be expressed uniquely as the sum of a symmetric and an anti-symmetric matrix.
@@
Therefore, only the symmetric part ${\bm{B}}_{a::}^{sym}$ contributes. From here on, we drop the $\cdot^{sym}$ superscript and assume the symmetric form for any interaction matrix or bilinear tensor ($b_{aij}=\frac{1}{2}(w_{ai}v_{aj}+w_{aj}v_{ai})$). Symmetric matrices have simpler eigendecompositions since the eigenvalues are all real-valued and the eigenvectors are orthogonal, by the spectral theorem.
Incorporating biases. If a bilinear layer has biases, we can augment the weight matrices to adapt our approach. Given activations of the form $g({\bm{x}})=(W{\bm{x}}+{\bm{b}}_{1})\odot({\bm{V}}{\bm{x}}+{\bm{b}}_{2})$, define ${\bm{W}}^{\prime}=[{\bm{W}};{\bm{b}}_{2}]$, ${\bm{V}}^{\prime}=[{\bm{V}};{\bm{b}}_{2}]$, and ${\bm{x}}^{\prime}=[{\bm{x}},1]$. Then, $g({\bm{x}})=({\bm{W}}^{\prime}{\bm{x}}^{\prime})\odot({\bm{V}}^{\prime}{\bm{x}}^{\prime})$ in a bilinear layer with biases. In subsection 4.3, we study a toy classification task using a model trained with biases, illustrating how biases can be interpreted using the same framework. For the rest of our experiments, we used models without biases for simplicity, as it did not harm performance. See Appendix J for details.
@@ 3 Analysis methods
Since bilinear MLPs can be expressed in terms of a third-order tensor, $\bm{\mathsfit{B}}$}, they can be flexibly analyzed using techniques from linear algebra, such as decompositions and transformations. The choice of analysis approach depends on what additional information, in terms of previously obtained input or output features, is provided.
@@ 3.1 Input / output features $\rightarrow$ direct interactions
If we have already obtained meaningful sets of features for the bilinear MLP’s inputs and outputs, for example, by training sparse autoencoders (SAEs), then we can directly study the interactions between these features. Using the feature dictionaries $D_{\text{in}}$ and $D_{\text{out}}$, we can transform the bilinear tensor into the feature basis via $\tilde{b}_{abc}=\sum_{ijk}d^{\text{\ out}}_{ai}\ b_{ijk}\ d^{\text{\ in}}_{bj}\ d^{\text{\ in}}_{ck}$. For a given set of sparse input and output activations, only a small subset of the interactions (with $a,b,c$ all active) will contribute, and the statistics of these active interactions can be studied.
If, instead, we want to determine the output feature activations from the inputs, we would use the output SAE’s encoder directions: $\tilde{b}_{abc}=\sum_{ijk}e^{\text{\ out}}_{ai}\ b_{ijk}\ d^{\text{\ in}}_{bj}\ d^{\text{\ in}}_{ck}$. In section 5, we use this approach to identify the top relevant interactions for features in a language model.
@@ 3.2 Output features $\rightarrow$ eigendecomposition
Given a set of meaningful features for the MLP outputs, we can identify the most important input directions that determine the output feature activations. The output features could come from a sparse autoencoder, from the unembedding layer (demonstrated in section 4), or from a previous decompilation of the later layers.
The interaction matrix, ${\bm{Q}}={\bm{u}}\cdot_{\text{out}}{\bm{\mathsfit{B}}}$ for a given output feature $\bm{u}$} can be decomposed into a set of eigenvectors (Figure 1). Since $\bm{Q}$} can be considered symmetric without loss of generality (see section 2), the spectral theorem gives
@@
with a set of $d$ (rank of $\bm{W}$}, $\bm{V}$}) orthonormal eigenvectors ${\bm{v}}_{i}$ and real-valued eigenvalues $\lambda_{i}$. In the eigenvector basis, the output in the $\bm{u}$}-direction is
@@
where each term can be considered the activation for the eigenvector ${\bm{v}}_{i}$. That is, the bilinear layer’s outputs are quadratic in the eigenvector basis.
@@ 3.3 No features $\rightarrow$ higher-order SVD
If we have no prior features available, it is still possible to determine the most important input and output directions of $\bm{\mathsfit{B}}$} through a higher-order singular value decomposition (HOSVD). The simplest approach that takes advantage of the symmetry in $\bm{\mathsfit{B}}$} is to reshape the tensor by flattening the two input dimensions to produce a $d_{\text{out}}\times d_{\text{in}}^{2}$ shaped matrix and then do a standard singular value decomposition (SVD). Schematically, this gives
@@
where $\bm{q}$} can still be treated as an interaction matrix and further decomposed into eigenvectors as described above.
@@ 4 Image classification: interpreting visual features
We consider models trained on the MNIST dataset of handwritten digits and the Fashion-MNIST dataset of clothing images. This is a semi-controlled environment that allows us to evaluate the interpretability of eigenvectors on a nontrivial task. This section analyses a shallow feedforward network (FFN) consisting of an embedding projection, a bilinear layer, and a classification head; see Appendix F for details.
@@ 4.1 Qualitative assessment: top eigenvectors appear interpretable
The eigenvectors are derived using the unembedding directions for the digits to obtain interaction matrices that are then decomposed following subsection 3.2. We can visualize them by projecting them into the input space using the embedding weights. Because the activation of an eigenvector $\bm{v}$} with eigenvalue $\lambda_{i}$ is quadratic in the input, $\lambda({\bm{v}}^{T}{\bm{x}})^{2}$, the sign of the eigenvector $\bm{v}$} is arbitrary. The quadratic leads to XOR-like behavior where high overlap with an eigenvector’s positive regions (blue) or the negative regions (red)—but not both—leads to large activation magnitude, while the overall sign is determined by the eigenvalue (Figure 2A).
For MNIST, the top positive eigenvector for each output class emphasizes a curve segment specific to its digit or otherwise resembles a prototypical class image (Figure 2B). Top eigenvectors for FMNIST function as edge detectors, focusing on important edges for each clothing article, such as the leg gap for trousers. Edge detecting relies on the XOR-like behavior of the eigenvector’s quadratic activation.
Figure 2: A) Eigenvector activations are quadratic in the input and have a large magnitude if an input aligns with the positive (blue) regions or the negative (red) regions, but not both. B) Top eigenvectors for single-layer MNIST and Fashion-MNIST models, revealing the most significant patterns learned for each class. In MNIST, eigenvectors represent components of the target class, while Fashion-MNIST eigenvectors function as edge detectors. Best viewed in color.
@@
Only a small fraction of eigenvalues have non-negligible magnitude (Figure 3). Different top eigenvectors capture semantically different aspects of the class. For example, in the spectrum for digit 5, the first two positive eigenvectors detect the 5’s horizontal top stroke but at different positions, similar to Gabor filters. The next two positive eigenvectors detect the bottom segment. The negative eigenvectors are somewhat less intuitive but generally correspond to features that indicate the digit is not a five, such as an upward curve in the top right quadrant instead of a horizontal stroke. In Appendix C, we study this technique towards explaining an input prediction. Details of the training setup are outlined in Appendix F while similar plots for other digits can be found in Appendix A.
Because we can extract features directly from model weights, we can identify overfitting in image models by visualizing the top eigenvectors and searching for spatial artifacts. For instance, the eigenvectors of unregularized models focus on certain outlying pixels (Figure 4). We found adding dense Gaussian noise to the inputs (Bricken et al., 2023a) to be an effective model regularizer, producing bilinear layers with more intuitively interpretable features. Increasing the scale of the added noise results produces more digit-like eigenvectors and results in a lower-rank eigenvalue spectrum (Appendix D). These results indicate that our technique can qualitatively help uncover overfitting or other unwanted behavior in models. Furthermore, it can be used to evaluate the effect of certain regularizers and augmentation techniques, as explored in Appendix B.
@@ 4.2 Quantitative assessment: eigenvectors learn consistent patterns
Both the ordering and contents of top eigenvectors are very consistent across runs. The cosine similarities of the top eigenvector are between 0.8 and 0.9 depending on size (Figure 5). Generally, increasing model sizes results in more similar top eigenvectors. Further, truncating all but the top few eigenvectors across model sizes yields very similar classification accuracy. This implies that, beyond being consistently similar, these eigenvectors have a comparable impact on classification. In Appendix E, we further study the similarity of eigenvectors between sizes and show that retaining only a handful of eigenvectors results in minimal accuracy drops (0.01%).
@@ 4.3 Comparing with ground truth: eigenvectors find computation
To perform a ground-truth assessment of eigenvectors, we consider a task from a mechanistic interpretability challenge, where the goal was to determine the labeling function (training objective) from a model Casper (2023). Specifically, the challenge required reverse-engineering a binary image classifier trained on MNIST, where the label is based on the similarity to a specific target image. The model predicted ‘True’ if the input has high cosine similarity to this target or high similarity to the complement of that target  [^1] By complement, we mean one minus the grayscale value of the target. and ‘False’ otherwise. This target is chosen as an instance of a ‘1’.
@@
We perform our decomposition on the output difference ($\text{True}-\text{False}$) since this is the only meaningful direction before the softmax. This consistently reveals one high positive eigenvalue; the rest are (close to) zero (Figure 6). The most positive eigenvector is sufficient for completing the task; it computes the exact similarity we want. If the input is close to the target, the blue region will match; if it is close to the complement, the red will match; if both are active simultaneously, they will somewhat cancel out. The remaining two eigenvectors are separated as they seem to overfit the data slightly; the negative eigenvector seems to penalize diagonal structures.
Contrary to other models, this task greatly benefited from including biases. This arises from the fact that the model must not only compute similarity but also make its binary decision based on a learned threshold. If no bias is provided, the model attempts to find quadratic invariances in the data, which don’t generalize well, especially given the important but sensitive role of this threshold in classification. Here, the bias (shown in the bottom corner of Figure 6) represents a negative contribution. The role of biases in bilinear layers is further discussed in Appendix J.

### v2

@@ 1 Introduction
One path to better circuit discovery is to use more inherently interpretable architectures. Previous work constrains the model into using human-understandable components or concepts (Koh et al., 2020; Chen et al., 2019), but this typically requires labeled training data for the predefined concepts and involves a trade-off in accuracy compared to learning the best concepts for performance (Henighan et al., 2023). Ideally, we could more easily extract and interpret the concepts that models naturally learn rather than force the model to use particular concepts. To this end, Sharkey (2023) suggested that bilinear layers Lin et al. (2015); Li et al. (2017); Chrysos et al. (2021) of the form $g({\bm{x}})=({\bm{W}}{\bm{x}})\odot({\bm{V}}{\bm{x}})$ are intrinsically interpretable because their computations can be expressed in terms of linear operations with a third order tensor. This enables the use of tensor or matrix decompositions to directly understand the weights. Moreover, bilinear layers outperform ReLU-based transformers in language modeling (Shazeer, 2020) and have performance only slightly below SwiGLU, which is prevalent in competitive models today (Touvron et al., 2023).
Tensor decompositions have long been studied in machine learning (Cichocki et al., 2015; Panagakis et al., 2021; Sidiropoulos et al., 2017) where most applications are based on an input dataset. Using decompositions to extract features directly from the weights of tensor-based models is less explored. Here, we show that bilinear MLPs can be decomposed into functionally relevant, interpretable components by directly decomposing the weights, without using inputs. These decompositions reveal a low-rank structure in bilinear MLPs trained across various tasks. In summary, this paper demonstrates that bilinear MLPs are an interpretable drop-in replacement for ordinary MLPs in a wide range of settings. Our contributions are as follows:
@@ 2 Background
[…]
Defining bilinear MLPs. Modern Transformers (Touvron et al., 2023) feature Gated Linear Units (GLUs), which offer a performance gain over standard MLPs for the same number of parameters (Shazeer, 2020; Dauphin et al., 2017). GLU activations consist of the component-wise product of two linear up-projections of size $(d_{\text{hidden}},d_{\text{input}})$, ${\bm{W}}$ and ${\bm{V}}$, one of which is passed through a nonlinear activation function $\sigma$(Equation 1). The hidden activations $g({\bm{x}})$ then pass through a down-projection ${\bm{P}}$ of size $(d_{\text{output}},d_{\text{hidden}})$. We omit biases for brevity.
@@
A bilinear layer is a GLU variant that omits the nonlinear activation function $\sigma$. Bilinear layers beat ordinary ReLU MLPs and perform almost as well as SwiGLU on language modeling tasks (Shazeer, 2020). We corroborate these findings in Appendix I, and show that bilinear layers achieve equal loss when keeping training time constant and marginally worse loss when keeping data constant.
@@
We call the $(d_{\text{input}},d_{\text{input}})$ matrix ${\bm{w}}_{a:}{\bm{v}}_{a:}^{T}={\bm{B}}_{a::}$ an interaction matrix since it defines how each pair of inputs interact for a given output dimension $a$.
[…]
Simplifications due to symmetry. Because an interaction matrix is always evaluated with two copies of the input ${\bm{x}}$, it contains redundant information that does not contribute to the activation. Any square matrix can be expressed uniquely as the sum of a symmetric and anti-symmetric matrix.
@@
Therefore, only the symmetric part ${\bm{B}}_{a::}^{sym}$ contributes. From here on, we drop the $\cdot^{sym}$ superscript and assume the symmetric form for any interaction matrix or bilinear tensor ($b_{aij}=\frac{1}{2}(w_{ai}v_{aj}+w_{aj}v_{ai})$). Symmetric matrices have simpler eigendecompositions since the eigenvalues are all real-valued, and the eigenvectors are orthogonal by the spectral theorem.
Incorporating biases. If a bilinear layer has biases, we can augment the weight matrices to adapt our approach. Given activations of the form $g({\bm{x}})=({\bm{W}}{\bm{x}}+{\bm{b}}_{1})\odot({\bm{V}}{\bm{x}}+{\bm{b}}_{2})$, define ${\bm{W}}^{\prime}=[{\bm{W}};{\bm{b}}_{2}]$, ${\bm{V}}^{\prime}=[{\bm{V}};{\bm{b}}_{2}]$, and ${\bm{x}}^{\prime}=[{\bm{x}},1]$. Then, $g({\bm{x}})=({\bm{W}}^{\prime}{\bm{x}}^{\prime})\odot({\bm{V}}^{\prime}{\bm{x}}^{\prime})$ in a bilinear layer with biases. In subsection 4.3, we study a toy classification task using a model trained with biases, illustrating how biases can be interpreted using the same framework. For the rest of our experiments, we used models without biases for simplicity, as it did not harm performance. See Appendix L for details.
@@ 3 Analysis methods
[…]
@@ 3.1 Input / output features $\rightarrow$ direct interactions
If we have already obtained meaningful sets of features for the bilinear MLP’s inputs and outputs, for example from a set of latent feature dictionaries $F^{\text{in}}$ and $F^{\text{out}}$, then we can directly study the interactions between these features and understand how the output features are constructed from input ones. We can transform the bilinear tensor into the feature basis via $\tilde{b}_{abc}=\sum_{ijk}f^{\text{\ out}}_{ai}\ b_{ijk}\ f^{\text{\ in}}_{bj}\ f^{\text{\ in}}_{ck}$. For a given set of sparse input and output activations, only a small subset of the interactions (with $a,b,c$ all active) will contribute, and the statistics of these active interactions can be studied.
For dictionaries obtained from sparse autoencoders (SAEs) we can instead use the output SAE’s encoder directions in the transformation: $\tilde{b}_{abc}=\sum_{ijk}e^{\text{\ out}}_{ai}\ b_{ijk}\ f^{\text{\ in}}_{bj}\ f^{\text{\ in}}_{ck}$. Then the output activations are $z^{\text{out}}_{c}=\text{ReLU}(\sum_{ab}\tilde{b}_{abc}z^{\text{in}}_{a}z^{\text{in}}_{b})$ in terms of the input directions ${\bm{z}}^{\text{in}}$. In section 5, we use this approach to identify the top relevant interactions for features in a language model.
@@ 3.2 Output features $\rightarrow$ eigendecomposition
Given a set of meaningful features for the MLP outputs, we can identify the most important input directions that determine the output feature activations. The output features could come from a dictionary, from the unembedding (shown in section 4), or from the decompilation of later layers.
[…]
@@
with a set of $d$ (the rank of ${\bm{W}}$, ${\bm{V}}$) orthonormal eigenvectors ${\bm{v}}_{i}$ and real-valued eigenvalues $\lambda_{i}$. In the eigenvector basis, the output in the ${\bm{u}}$-direction is
@@
where each term can be considered the activation for the eigenvector ${\bm{v}}_{i}$ of size $(d_{\text{input}})$. That is, the bilinear layer’s outputs are quadratic in the eigenvector basis.
@@ 3.3 No features $\rightarrow$ higher-order SVD
[…]
@@
where ${\bm{q}}$ can still be treated as an interaction matrix and further decomposed into eigenvectors as described above. We demonstrate this approach for an MNIST model in Appendix D.
@@ 4 Image classification: interpreting visual features
We consider models trained on the MNIST dataset of handwritten digits and the Fashion-MNIST dataset of clothing images. This is a semi-controlled environment that allows us to evaluate the interpretability of eigenvectors computed using the methods in subsection 3.2. This section analyses a shallow feedforward network (FFN) consisting of an embedding projection, a bilinear layer, and a classification head; see Appendix G for details.
@@ 4.1 Qualitative assessment: top eigenvectors appear interpretable
[…]
We can visualize them by projecting them into the input space using the embedding weights. Because the activation of an eigenvector ${\bm{v}}$ with eigenvalue $\lambda_{i}$ is quadratic in the input, $\lambda({\bm{v}}^{T}{\bm{x}})^{2}$, the sign of the eigenvector ${\bm{v}}$ is arbitrary. The quadratic leads to XOR-like behavior where high overlap with an eigenvector’s positive regions (blue) or the negative regions (red)—but not both—leads to large activation magnitude, while the overall sign is determined by the eigenvalue (Figure 2A).
[…]
Figure 2: A) Eigenvector activations are quadratic in the input and have a large magnitude if an input aligns with the positive (blue) regions or the negative (red) regions, but not both. B) Top eigenvectors for single-layer MNIST and Fashion-MNIST models, revealing the most significant patterns learned for each class. In MNIST, eigenvectors represent components of the target class, while Fashion-MNIST eigenvectors function as localized edge detectors. Best viewed in color.
@@
For MNIST, the top positive eigenvector for each output class emphasizes a curve segment specific to its digit or otherwise resembles a prototypical class image (Figure 2B). Top eigenvectors for FMNIST function as localized edge detectors, focusing on important edges for each clothing article, such as the leg gap for trousers. The localized edge detection relies on the XOR-like behavior of the eigenvector’s quadratic activation.
Only a small fraction of eigenvalues have non-negligible magnitude (Figure 3). Different top eigenvectors capture semantically different aspects of the class. For example, in the spectrum for digit 5, the first two positive eigenvectors detect the 5’s horizontal top stroke but at different positions, similar to Gabor filters. The next two positive eigenvectors detect the bottom segment. The negative eigenvectors are somewhat less intuitive but generally correspond to features that indicate the digit is not a five, such as an upward curve in the top right quadrant instead of a horizontal stroke. In Appendix C, we study this technique towards explaining an input prediction. Details of the training setup are outlined in Appendix G while similar plots for other digits can be found in Appendix A.
Because we can extract features directly from model weights, we can identify overfitting in image models by visualizing the top eigenvectors and searching for spatial artifacts. For instance, the eigenvectors of unregularized models focus on certain outlying pixels (Figure 4). We found adding dense Gaussian noise to the inputs (Bricken et al., 2023a) to be an effective model regularizer, producing bilinear layers with more intuitively interpretable features. Increasing the scale of the added noise results produces more digit-like eigenvectors and results in a lower-rank eigenvalue spectrum (Appendix E). These results indicate that our technique can qualitatively help uncover overfitting or other unwanted behavior in models. Furthermore, it can be used to evaluate the effect of certain regularizers and augmentation techniques, as explored in Appendix B.
@@ 4.2 Quantitative assessment: eigenvectors learn consistent patterns
Both the ordering and contents of top eigenvectors are very consistent across runs. The cosine similarities of the top eigenvector are between 0.8 and 0.9 depending on size (Figure 5). Generally, increasing model sizes results in more similar top eigenvectors. Further, truncating all but the top few eigenvectors across model sizes yields very similar classification accuracy. This implies that, beyond being consistently similar, these eigenvectors have a comparable impact on classification. In Appendix F, we further study the similarity of eigenvectors between sizes and show that retaining only a handful of eigenvectors results in minimal accuracy drops (0.01%).
@@ 4.3 Comparing with ground truth: eigenvectors find computation
To perform a ground-truth assessment of eigenvectors, we consider a task from a mechanistic interpretability challenge, where the goal was to determine the labeling function (training objective) from a model Casper (2023). Specifically, the challenge required reverse-engineering a binary image classifier trained on MNIST, where the label is based on the similarity to a specific target image. The model predicted ‘True’ if the input has high cosine similarity to this target or high similarity to the complement (one minus the grayscale) of that target and ‘False’ otherwise. This target is chosen as an instance of a ‘1’.
@@
[…]
Contrary to other models, this task greatly benefited from including biases. This arises from the fact that the model must not only compute similarity but also make its binary decision based on a learned threshold. If no bias is provided, the model attempts to find quadratic invariances in the data, which don’t generalize well, especially given the important but sensitive role of this threshold in classification. Here, the bias (shown in the bottom corner of Figure 6) represents a negative contribution. The role of biases in bilinear layers is further discussed in Appendix L.